OpenStudy (anonymous):

given the verticeS P(-5,2), Q(-1,-5), R(9,10) prove uSing Slopes that PQR is a right triangle

OpenStudy (anonymous):

Find triangle's sides and they should be: a<b+c b<a+c c<a+b

OpenStudy (asnaseer):

find the slopes: 1) between P and Q 2) between Q and R 3) between P and R if two lines are perpendicular to each other, then the product of their slopes = -1 use this fact to show PQR is a right triangle.

OpenStudy (anonymous):

there are many ways to skin a cat

OpenStudy (anonymous):

is slope is same word as side?

OpenStudy (asnaseer):

no

OpenStudy (asnaseer):

slope = rise/run

OpenStudy (anonymous):

Things get tricky when you find that one of the slopes is NaN

OpenStudy (anonymous):

 so PR PQ RQ is sides and slopes isn't it?

OpenStudy (anonymous):

well, each side has a slope associated with it :-P

OpenStudy (asnaseer):

for two point with coordinate (x1, y1) and (x2, y2), slope = (y2-y1)/(x2-x1) side = distance between these two points

OpenStudy (anonymous):

well what do you get ?

OpenStudy (anonymous):

\(y=\mbox{k}x+\mbox{b}\) k is slope right?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

@mikow - the 3 slopes I get are \(\frac{-7}{4}, \frac{3}{2} and \frac{4}{7}\)

OpenStudy (anonymous):

the products of the slopes -7/4 and 4/7 is -1, so the two lines with those slopes denote the right angle of the right triangle

OpenStudy (anonymous):

@asnaseer .. we're the Same :)

OpenStudy (anonymous):

what about poor @agdgdgdgwngo :(

OpenStudy (asnaseer):

good - glad you got there in the end :-) @agdg... (why is your name tag so difficult to type) - yes you were also right :-)

OpenStudy (anonymous):

dont be Sad :)

OpenStudy (anonymous):

be happy like JOLLIBEE AHAHAHHA LOL

OpenStudy (anonymous):

it;s easy to type:: 1. place your left pinky on the shift button, your left ring-finger on the a button, and your index finger on the g button. 2. press shift and a 3. release shift 4. press g 5. move your middle finger over the d button. 6 repeat three times: press the d button and the g button, in order 7. type in "wngo"

OpenStudy (asnaseer):

LOL! - do you hold special training sessions that I can attend? ;-)

OpenStudy (anonymous):

yeah only 50 (insert your currency) per hour!

OpenStudy (anonymous):

tnx to both of you :)

OpenStudy (asnaseer):

@agdgdgdgwngo - JUST in case case your name really is agdgdgdgwngo - I mean no offence by my remarks - it's all in jest. and I can only type with my index fingers as I never mastered touch typing :-(

OpenStudy (anonymous):

no shame in hunt-and-peck style typing : http://en.wikipedia.org/wiki/Typing#Hunt_and_peck

OpenStudy (anonymous):

lol^

OpenStudy (asnaseer):

Wow! I never new there was a term for my disability ;-)

OpenStudy (anonymous):

wait, its a disablility

OpenStudy (asnaseer):

:-D

OpenStudy (asnaseer):

lets just call it a "special" talent shall we :-)

OpenStudy (anonymous):

 show that the nearest distance from the common point to the plane is lA.BxCl / lAxB+BxC+CxAl­­